OpenStudy (anonymous):

Is the triangle formed by sides with lengths 5, 12, and 13 a right triangle? a. true b. false

OpenStudy (anonymous):

What have you got so far?

OpenStudy (anonymous):

true

OpenStudy (anonymous):

lol, you just stated an answer where I was looking for a process, do you know why it is true or false?

OpenStudy (anonymous):

if h^2=b^2+p^2 then its a right triangel

OpenStudy (anonymous):

13^2=5^2+12^2 169=169 its true

OpenStudy (anonymous):

yup...5^=4^2+3^2